OpenStudy (anonymous):

Show work: x/(x+1) < x/(x-1) what I have so far: x/(x+1) - x/(x-1)(x+1)(x-1)= 0(x+1)(x-1)... because you have to multiply both sides by the LCM which gives me: x(x-1) - x(x+1) = 0 then x^2 - 1x - x^2 + 1x=0 which is 0 = 0 which doesnt work! what did i do wrong?

OpenStudy (anonymous):

oops sorry, it should be x(x-1) - x(x+1) = 0 then x^2-1x-x^2-1x= 0 which is -2x=0 which is still 0 = 0

OpenStudy (anonymous):

\[\frac{x}{(x+1)} < \frac{x}{(x-1)}\] then: \[\frac{x}{(x+1)} - \frac{x}{(x-1)}<0\] then: \[\frac{x(x-1)-x(x+1)}{(x+1)(x-1)}<0\] then: \[\frac{x^2-x-x^2-x}{(x+1)(x-1)}<0\] then: \[\frac{-2x}{(x+1)(x-1)}<0\] Now you have to find the sign of the numerator and the sign of the denominator. Then superpose them. Please, tell me if you know how to continue.

OpenStudy (anonymous):

shouldn't it be the equation > 0?

OpenStudy (anonymous):

and cant i get rid of the (x+1) and (x-1) in the denominator when i multiply the equation by (x+1)(x-1) which is the LCM?

OpenStudy (anonymous):

No, because on the first step we substract both sides x/(x-1). And substracting doesn't make change in the sign.

OpenStudy (anonymous):

No, you can't get rid of that when you do the LCM.

hero (hero):

Alternatively, you can perform these steps: 1. Note that \(x \ne 1\) \(\large\frac{x}{x+1} < \frac{x}{x-1}\) 2. Cross Multiply: \(x(x-1) < x(x+1)\) 3. Distribute: \(x^2 - x < x^2 + x\) 4. Subtract x^2 from both sides: \(-x < + x\) 5. Add x to both sides \(0 < 2x\) 6. Divide both sides by 2 \(x > 0\) So we know that x > 0, except where x = 1

OpenStudy (anonymous):

@Hero We can't do the 2nd step: We have to check that x+1 and x-1 are >0, Because if any of them are <0 we have to change the sign of the inequality.

hero (hero):

once we solve for x, we'll know if it is greater than zero or not. Algebraically, cross-multiplication is a legal step.

hero (hero):

In this case, x will not and cannot be less than or equal to zero

OpenStudy (anonymous):

the answer is -1<x<0 and x>-1 i think

OpenStudy (anonymous):

@Hero I do not agree with that. Here you can see that what you are saying is wrong: http://www.wolframalpha.com/input/?i=x%2F%28x%2B1%29+%3C+x%2F%28x-1%29 The solution is x>1 AND -1<x<0

hero (hero):

Yeah, I neglected to include the part where -1<x<0

OpenStudy (anonymous):

oops correction to my comment -1<x<0 and x<1

OpenStudy (anonymous):

Thanks guys! :)

OpenStudy (anonymous):

@Elysse2012 Is there any doubt? Be careful while working with this kind of problems. Be careful with the "cross multiply" I hope it helps! :D

OpenStudy (anonymous):

No doubt anymore! Thanks to the both of you!

OpenStudy (anonymous):

You deserve that! Thanks for the medal! :D

hero (hero):

My explanation was actually incomplete. If @radabc02, had not interrupted my explanation, I would have came to the same conclusion

hero (hero):

What you do is note that \(x \ne -1\) and \(x \ne 1\). The zero is another boundary that should be noted. Doing so would give the boundaries:  Which after testing, we see that -1<x<0 and x>1